# Interactions across multiple games: cooperation, corruption, and organizational design

### v2

@@ 4 Assigning Teams to Tasks
The more interesting case is one where the full-specialization team structure does not yield such optimal cooperation; therefore, we investigate situations in which sustaining cooperation in the good game can only be done if it is mixed with some of the bad game ($d_{G}>\frac{\delta}{1-\delta}a_{G}c_{G}$), and there is cooperation on both games when teams face a high enough probability of bad tasks ($\max_{g}\{d_{g}\}<\frac{\delta}{1-\delta}a_{B}c_{B}$). In these situations, the optimal team structure may involve creating two types of team assignments: one that includes the bad task with the minimum share needed to sustain total cooperation, and the other type of team assignment covers the remainder to ensure that all tasks are covered—which are either all good or all bad, depending on parameters—according to the exogenous constraint given in (3).
Thus, consider assignment $\nu^{coop}$ that places just enough weight on the bad task to incentivize cooperation in both games, i.e.,
@@
Either the above assignment is part of an optimal team structure, or the social cost of bad cooperation is too large and no cooperation is optimal. In the former case, the share of the population with assignment $\nu^{coop}$ is maximized so that on the remaining assignment, teams face only good or only bad tasks and are constantly reshuffled to preclude cooperation. The next proposition formalizes this.
@@
  • $\nu^{g}$ with $\nu^{g}(t_{g})=1$, for some $g\in\{G,B\}$, and $r_{\nu^{g}}$ sufficiently high to preclude cooperation,
@@
Figure 2 depicts case (i) for two subcases. In Panel (a), total cooperation is possible if there are only generalists, i.e., game probabilities are given by $a_{G}q_{G},a_{B}q_{B}$, captured by point $x^{gen}$. However, the designer can improve the social value by assigning some teams to relatively more good tasks; the resulting task assignment corresponds to point $x^{coop}$. Those teams are not reshuffled. The remaining teams are assigned only to the bad task and then reshuffled; if they were not reshuffled, they would cooperate on bad games ($x^{spec}_{\text{fixed}}$); however, as they are frequently reshuffled, they do not cooperate ($x^{spec}_{\text{reshuffled}}$). In Panel (b), total cooperation is not sustainable with only generalists. There, the teams that can support cooperation are the ones assigned to more bad tasks, and the specialized assignment is to good tasks only to satisfy the balance condition (3).
@@
This discussion suggests that organizations may benefit by specializing their teams and reshuffling them at different rates. In the example of police officers being tasked with either responding to emergencies or traffic policing, good cooperation among emergency responders—backing up one’s partner in dangerous situations—may be supported by keeping officers together, and bad cooperation—colluding in taking bribes—can be stopped by reshuffling partnerships often.
@@ 5 Dynamic, Reactive Assignments
In Section 4, teams could be (randomly) assigned to different tasks; in this section, we enrich the analysis to allow such assignments to be dynamic (i.e., changing over time) and possibly reactive (i.e., depending on past assignments and realized games).
We begin with situations in which the designer can observe whether a good game has arrived (but not which actions were played). For example, the designer can observe whether a team in the army engaged with the enemy but not how they cooperated internally. It may be natural to also assume that the designer does not observe whether a bad game—e.g., an opportunity to take a bribe—arrived; this turns out to be inconsequential in our results, as we show. After first treating the case where the arrival of a good game is observed, we also consider the case where the designer cannot observe whether a good game actually arrived in a given period or not.
To illustrate how a dynamic, reactive assignment can strictly improve the value obtained by the designer compared to a constant (random) assignment, consider the following example.
Example 1.
Suppose that if the designer cannot react and must use a constant (random) assignment as in Section 4, then a probability of $2/3$ on the bad task is needed to incentivize cooperation in both games (i.e., $\nu^{coop}(t_{G})=1/3$ and $\nu^{coop}(t_{B})=2/3$). [^9] For instance, let $\delta=\frac{3}{5},c_{G}=c_{B}=1,a_{G}=\frac{1}{2},a_{B}=\frac{3}{4},d_{G}=1$, and $d_{B}=\frac{1}{2}$. Then, (4) is satisfied as $\max_{g}\{d_{g}\}=\max\{\frac{1}{2},1\}=\frac{\frac{3}{5}}{1-\frac{3}{5}}(\frac{1}{3}\cdot\frac{1}{2}\cdot 1+\frac{2}{3}\cdot\frac{3}{4}\cdot 1)=\frac{\delta}{1-\delta}\left(\nu^{coop}(t_{G})a_{G}c_{G}+(1-\nu^{coop}(t_{G}))a_{B}c_{B}\right)$. Since we maintain $d_{G}>\frac{\delta}{1-\delta}a_{G}c_{G}$ and $\max_{g}\{d_{g}\}<\frac{\delta}{1-\delta}a_{B}c_{B}$, it follows that $a_{B}c_{B}>a_{G}c_{G}$ and so a future bad task provides stronger incentives to cooperate today than a future good task.
@@
Thus, the good task followed immediately by two bad tasks—i.e., $t_{G},t_{B},t_{B},t_{G},t_{B},t_{B},t_{G},t_{B},t_{B},\dots$—provides stronger incentives to cooperate on good games than the random continuation. In particular, if $d_{G}>d_{B}$, and so given $\nu^{coop}$, players are indifferent between cooperating or not when facing a good game. If cooperation in the bad games is still strictly incentivized (i.e., $d_{B}$ is small enough), players are now strictly incentivized to cooperate both after good and bad games. [^10] Given the parameters in footnote 9, players are still strictly incentivized to cooperate after a bad game as $d_{B}=\frac{1}{2}<\frac{\frac{3}{5}}{1-\frac{3}{5}}\frac{1}{2}\cdot 1=\frac{\delta}{1-\delta}a_{G}c_{G}$. That is, even only future good games would suffice to ensure cooperation on current bad games. The designer can use this slackness in incentives to use fewer bad tasks on average. Instead of following a good task with a bad task, the designer follows a good task with one bad task for sure and then a probability of a bad task below one (with the remaining probability returning to the good task early). This uses fewer bad tasks, and the remaining bad tasks can be assigned to some constantly reshuffled team to avoid cooperation in the associated bad games. Overall, this is a strict improvement for the designer.
@@
An illustration of a reactive assignment like that described in Example 1 is immediately assigning a military unit to noncombat status after it has faced combat and then, at some further time, reassigning those soldiers back to the front where combat becomes more likely again.
Next, we formalize this setup. We then characterize when the designer benefits from dynamically and reactively assigning teams to tasks, as in Example 1. Having established when the designer benefits, we consider how and show that in a designer-optimal reactive team structure, teams can be grouped into two types: those who always (on-path) cooperate and those who do not. To further characterize the optimal solution, we consider the most relevant case in which the incentive to deviate in bad games, $d_{B}$, is sufficiently low so that the incentives provided to cooperate in good tasks are the binding constraint. [^11] When $d_{G}$ is sufficiently low, full specialization ensures optimal cooperation—Observation 1. We then fully characterize the designer-optimal reactive team structure in Proposition 4. In an optimal structure, teams that cooperate are assigned to the good task until a good game arrives, in which case they are moved to the bad task for some amount of time before returning to the good task. Teams that never cooperate pick up the slack so that the exogenous task constraint is satisfied.
@@ 5.1 Definitions and a Preliminary Result
A reactive assignment is a function $m:(\mathcal{T}\times\{G,B,\emptyset\})^{K}\rightarrow\Delta(\mathcal{T})$, where $K$ is a positive integer.
@@
That is, instead of simply being a point in $\Delta(\mathcal{T})$, the function $m$ can condition on which previous tasks were assigned and which games, if any, were played up through the previous $K$ periods, and then (possibly randomly) assign a task. Let $\mathcal{M}$ denote the set of possible reactive assignments.
@@
A reactive assignment results in a distribution over tasks, which, when aggregated across a continuum of teams, appropriately staggered, leads to a steady-state frequency of tasks. Specifically, the function $m$ can also be viewed as a map from $(\mathcal{T}\times\{G,B,\emptyset\})^{K}$ into $\Delta\left((\mathcal{T}\times\{G,B,\emptyset\})^{K}\right)$, as it drops one entry and adds a new one each period. For $s,s^{\prime}\in\mathcal{S}\equiv(\mathcal{T}\times\{G,B,\emptyset\})^{K}$, let $m(s|s^{\prime})$ be the probability of transitioning from state $s^{\prime}$ to $s$. Then $m$ defines a Markov chain on $\mathcal{S}$.
A steady-state distribution induced by $m$ is denoted by $\overline{m}$ and satisfies $\overline{m}(s)=\sum_{s^{\prime}\in\mathcal{S}}m(s|s^{\prime})\overline{m}(s^{\prime})$. $m$ could be periodic and reducible and hence may have multiple steady states.
@@
An organization’s reactive team structure is $\lambda\in\Delta(\mathcal{M})$, together with reshuffling rate $r_{m}$ and a steady-state distribution $\overline{m}$ for each $m\in{\mathrm{supp}}(\lambda)$. [^12] Note that the organization can maintain the steady state by assigning a $\overline{m}$ mass of teams to states in $m$ for each $m$. For example, in the assignment that alternates between good and bad tasks, even though a single team never reaches a steady state but instead keeps oscillating, a population of teams that are staggered would have half on $t_{G}$ and half on $t_{B}$ at any point in time. We thus assume the organization appropriately staggers teams to maintain the steady state. The history of a team and players’ strategies, specified for each $m$, are as in the previous section. Furthermore, we continue to model reshuffling as a reduction in the discount factor as, given the continuum of agents and a finite set of assignments and states, players can be matched to a new team in the same reactive assignment and state.
@@
Lastly, the dynamic team structure must respect a task-assignment constraint—the equivalent of (3). Let $\mathcal{S}^{t}(m,\overline{m})\subseteq\mathcal{S}$ denote the set of states $s=[(t_{\tau},g_{\tau})]_{\tau=0}^{K}\in\mathcal{S}$ such that $t_{K}=t$ for $t\in\mathcal{T}$, i.e., those where the current task is task $t$. The overall constraint on a team structure is:
@@
The next lemma establishes that Properties 16 continue to imply a simple on-path equilibrium structure in which players only fully cooperate or fully not cooperate. As a result, there is a well-defined social value given any reactive team structure.
@@
$\sum_{m\in{\mathrm{supp}}(\lambda)}\lambda(m)\left(\sum_{s\in\mathcal{G}(m,r_{m})}\overline{m}(s)V_{G}-\sum_{s\in\mathcal{B}(m,r_{m})}\overline{m}(s)V_{B}\right).$ (6)

### v3

@@ 4 Assigning Teams to Tasks
The more interesting case is one where the full-specialization team structure does not yield such optimal cooperation; therefore, we investigate situations in which sustaining cooperation in the good game can only be done if there is a high-enough probability of future bad games ($d_{G}>\frac{\delta}{1-\delta}a_{G}c_{G}$ and $\max_{g}\{d_{g}\}<\frac{\delta}{1-\delta}a_{B}c_{B}$). In these situations, the optimal team structure may involve creating two types of team assignments: one that includes the bad task with the minimum share needed to sustain total cooperation (as in (3) in Example 1), and another that ensures all tasks are covered according to the exogenous constraint given in (4).
Thus, consider assignment $\nu^{coop}$, which generalizes (3) in Example 1, that places just enough weight on the bad task to incentivize cooperation in both games, i.e.,
@@
Either assignment $\nu^{coop}$ is part of an optimal team structure, or the social cost of bad cooperation is too large and no cooperation is optimal. In the former case, the share of the population with assignment $\nu^{coop}$ is maximized, so that, under the assignment, teams face only good or only bad tasks and are constantly reshuffled to preclude cooperation. The next proposition formalizes this.
@@
  • $\nu^{g}$ with $\nu^{g}(t_{g})=1$, for some $g\in\{G,B\}$, and $r_{g}$ sufficiently high to preclude cooperation if g=Bg=B, [^10] If $g=G$, then there is no cooperation for any reshuffling rate since we are considering settings in which bad tasks are needed to sustain cooperation in good games.
@@
Let us revisit Example 1 and explicitly derive the optimal team structure for these parameters.
Continuation of Example 1.
A team structure $\lambda$ with positive weight on $\nu^{coop}$ can be optimal only if the social value coming from $\nu^{coop}$ is positive, i.e., if $\nu^{coop}(t_{G})a_{G}V_{G}+\nu^{coop}(t_{B})a_{B}V_{B}\geq 0$. In Example 1, we derived $\nu^{coop}$ to be $\nu^{coop}(t_{G})=1/3$ and $\nu^{coop}(t_{B})=2/3$. Since $a_{G}=1/2$ and $a_{B}=3/4$, we are in case (i) of the preceding proposition if $\frac{1}{3}\cdot\frac{1}{2}\cdot V_{G}+\frac{2}{3}\cdot\frac{3}{4}\cdot V_{B}>0$, i.e., $V_{G}>-3V_{B}$. (And if $V_{G}<-3V_{B}$, then any optimal team structure precludes all cooperation; if $V_{G}=-3V_{B}$, then optimal team structures have weight on $\nu^{coop}$ or on assignments that preclude cooperation.)
Suppose $V_{G}>-3V_{B}$ and so Proposition 2 implies that optimal team structure $\lambda$ puts maximal weight on $\nu^{coop}$ subject to satisfying the exogenous constraint on task coverage (4). Recall that half the teams need to be assigned to the good task and half to the bad task, i.e., $q_{G}=q_{B}=1/2$. Hence, to maintain task coverage, the designer assigns some teams to only good tasks, $\nu^{G}$; Panel (a) of Figure 2 depicts the relevant assignments. The relative weights of $\nu^{coop}$ and $\nu^{G}$ are found by noting that to cover all tasks, $\lambda$ must satisfy $\lambda(\nu^{coop})\nu^{coop}(t_{B})=q_{B}$, giving $\lambda(\nu^{coop})=3/4$ (which implies $\lambda(\nu^{G})=1/4$; and there is no cooperation for any $r_{G}$).
For other task coverage constraints, the optimal team structure changes. For instance, suppose that at any given point in time, only 25% of teams need to be assigned to the good task and now 75% to the bad task, i.e., $q_{G}=1/4$ and $q_{B}=3/4$. Again, the designer assigns some teams to $\nu^{coop}$, reducing the share of bad tasks relative to the generalist assignment, and others now to $\nu^{B}$ to maintain task coverage; Panel (b) of Figure 2 illustrates. The optimal team structure then satisfies $\lambda(\nu^{coop})\nu^{G}(t_{G})=q_{G}$, giving $\lambda(\nu^{coop})=3/4$ (which implies $\lambda(\nu^{B})=1/4$; $r_{\nu^{B}}$ is sufficiently high to preclude cooperation). [^12] To preclude cooperation on $\nu^{B}$, we require $d_{B}>\frac{(1-r_{\nu^{B}})\delta}{1-(1-r_{\nu^{B}})\delta}a_{B}c_{B}$, i.e., $\frac{1}{2}>\frac{(1-r_{\nu^{B}})\frac{3}{4}}{1-(1-r_{\nu^{B}})\frac{3}{4}}\cdot\frac{3}{4}\cdot 1,$ which simplifies to $r_{\nu^{B}}>\frac{7}{15}$.
@@
This discussion suggests that organizations may benefit by specializing their teams and reshuffling them at different rates. In the case of police officers tasked with either responding to emergencies or traffic policing, good cooperation among emergency responders—backing up one’s partner in dangerous situations—may be supported by keeping officers together, and bad cooperation—colluding in taking bribes—can be stopped by reshuffling partnerships often.
@@ 5 Dynamic, Reactive Assignments
In Section 4, teams could be (randomly) assigned to different tasks, but the assignment probabilities were the same in each period. We now further enrich the analysis to allow the designer to use assignments that change over time and possibly react to the past arrival process of games.
As we show, a designer can improve incentives and subsequently the social value by using a time-varying assignment strategy. The basic idea is as follows. Suppose that to provide incentives to cooperate in good games, a team needs to expect to see a certain amount of bad games in the future. That is, they get more motivation from the anticipation of bad game arrivals than from good ones. Under a random mixture, the possibility of seeing good games dampens the incentives from bad games. If, instead, once a team faces a good game, the designer promises to switch the team immediately to a task with only bad games for some time, then incentives to cooperate improve compared to having the team continue facing the mixture.
To illustrate this idea in more detail, consider again the parameterization in Example 1.
Continuation of Example 1.
Recall that $a_{G}=1/2,a_{B}=3/4$ and $c_{G}=c_{B}=1$. Hence, $a_{B}c_{B}>a_{G}c_{G}$, and so a future bad task provides stronger incentives to cooperate today than a future good task. (This is a general implication of our maintained assumption that $d_{G}>\frac{\delta}{1-\delta}a_{G}c_{G}$ and $\max_{g}\{d_{g}\}<\frac{\delta}{1-\delta}a_{B}c_{B}$.)
The designer can use this fact to maximize incentives when they are most needed. Specifically, instead of using a fixed random assignment of $\nu^{coop}$ with $\nu^{coop}(t_{G})=1/3$ and $\nu^{coop}(t_{B})=2/3$, suppose the designer uses a carefully orchestrated sequential assignment, so that the good task is always followed by two bad tasks and this then repeats—i.e., $t_{G},t_{B},t_{B},t_{G},t_{B},t_{B},t_{G},t_{B},t_{B},\dots$. Crucially, this provides stronger incentives to cooperate on good games, while the same fraction of bad tasks is covered. If, in addition, $d_{G}>d_{B}$, and so given $\nu^{coop}$, players are indifferent between cooperating or not when facing a good game, and cooperation in the bad games is still strictly incentivized (i.e., $d_{B}$ is small enough), then players are now strictly incentivized to cooperate both after good and bad games.
In the example presented here, players are, in fact, strictly incentivized to cooperate after a bad game since $d_{B}=\frac{1}{2}<\frac{\frac{3}{5}}{1-\frac{3}{5}}\frac{1}{2}\cdot 1=\frac{\delta}{1-\delta}a_{G}c_{G}$. That is, even being assigned to only good tasks (or any mixed assignment of tasks) in the future would suffice to ensure cooperation in the current bad game.
@@
The designer can now use this slack in incentives to reduce the average number of bad tasks in this assignment, and, with it, the overall share of bad games with cooperation. Instead of following a good task with two bad tasks, the designer follows a good task with one bad task for sure, and then a probability of a bad task below one (with the remaining probability returning to the good task early). This uses fewer bad tasks, and the remaining bad tasks can be assigned to some constantly reshuffled team to avoid cooperation in the associated bad games and to maintain task coverage. Overall, this is a strict improvement for the designer.
@@
An application of a reactive assignment like that described in the example above is immediately assigning a military unit to noncombat status after it has faced combat and, after some time, reassigning those soldiers back to the front where combat becomes more likely again.
The analysis here is a first example of dynamically adaptive mechanism design, in which a designer changes the mechanism agents play over time in response to past events. We now formalize the benefits of this additional assignment flexibility and characterize when the designer gains from dynamically and reactively assigning teams to tasks, as in the running example. We then show that in a designer-optimal reactive team structure, teams can be grouped into two types—those that always (on-path) cooperate and those that do not—and characterize each type’s assignments.
To formalize the ideas, we need to be explicit about what the designer can observe and condition their assignments on. To start, consider a situation in which the designer can observe whether a good game has arrived (but not which actions were played). For example, the designer can observe whether a team in the army engaged with the enemy, but not how they cooperated internally. It may be natural to also assume that the designer does not observe whether a bad game—e.g., an opportunity to take a bribe—arrived. Observability of bad game arrivals turns out to be inconsequential in our results. After first treating the case where the arrival of a good game is observed, we also consider the case where the designer cannot observe whether a good game actually arrived in a given period.
@@ 5.1 Definitions and a Preliminary Result
A reactive assignment is a function $m:(\mathcal{T}\times\{G,B,\emptyset\})^{K}\rightarrow\Delta(\mathcal{T})$, where $K$ is a positive integer. That is, instead of simply being a point in $\Delta(\mathcal{T})$, the function $m$ can condition on which previous tasks were assigned and which games, if any, were played up through the previous $K$ periods, and then (possibly randomly) assign a task. Let $\mathcal{M}$ denote the set of possible reactive assignments.
@@
A reactive assignment results in a distribution over tasks that, when aggregated across a continuum of teams, appropriately staggered, leads to a steady-state frequency of tasks. Specifically, the function $m$ can also be viewed as a map from $(\mathcal{T}\times\{G,B,\emptyset\})^{K}$ into $\Delta\left((\mathcal{T}\times\{G,B,\emptyset\})^{K}\right)$, as it drops one entry and adds a new one each period. For $s,s^{\prime}\in\mathcal{S}\equiv(\mathcal{T}\times\{G,B,\emptyset\})^{K}$, let $m(s|s^{\prime})$ be the probability of transitioning from state $s^{\prime}$ to $s$. Then $m$ defines a Markov chain on $\mathcal{S}$. A steady-state distribution induced by $m$ is denoted by $\overline{m}$ and satisfies $\overline{m}(s)=\sum_{s^{\prime}\in\mathcal{S}}m(s|s^{\prime})\overline{m}(s^{\prime})$. Note that since the Markov chain may be periodic or reducible—and thus admit multiple steady states—it is important to specify the relevant steady state.
To illustrate these definitions, consider again the assignment in Example 1 in which the good task is followed by one bad task for sure and then a bad task with probability $x$ (and with the remaining probability, 1x1-x, a return to the good task). In steady state, the process induced by $m$ visits only a small number of aggregate states (equivalence classes of histories that are relevant for continuation play under this assignment). In this example, we can summarize the relevant aggregate states for the designer to track by (i) the current task, (ii) whether a game arrives in the current period, and (iii) whether the team is in the first or second bad-task position of the cycle. This yields six aggregate states: $(t_{G},G)$, $(t_{G},\emptyset)$, $(t_{B},B)_{1}$, $(t_{B},\emptyset)_{1}$, $(t_{B},B)_{2}$, and $(t_{B},\emptyset)_{2}$. We thus consider the induced Markov chain on these aggregates, since all histories within a given aggregate generate the same continuation distribution under this assignment. The associated transition matrix is
[…]
For example, following a good task, teams move deterministically to the first bad task. Conditional on this transition, a bad game arrives with probability $a_{B}$, in which case the state becomes $(t_{B},B)_{1}$; if no game arrives, the state is $(t_{B},\emptyset)_{1}$. Then, teams transition to another bad task with probability $x$, and conditional on this transition, a bad game arrives with probability $a_{B}$, in which case the state becomes $(t_{B},B)_{2}$, and with complementary probability, the state becomes $(t_{B},\emptyset)_{2}$. With probability 1x1-x, teams transition back to the good task, and similarly either to state $(t_{G},G)$ or $(t_{G},\emptyset)$, etc. The resulting steady-state distribution over the six states is
$\overline{m}=\frac{1}{2+x}\Big(a_{G},\;1-a_{G},\;a_{B},\;1-a_{B},\;xa_{B},\;x(1-a_{B})\Big).$ (7)
@@
An organization’s reactive team structure is $\lambda\in\Delta(\mathcal{M})$, together with reshuffling rate $r_{m}$ and a steady-state distribution $\overline{m}$ for each $m\in{\mathrm{supp}}(\lambda)$. The history of a team and players’ strategies, specified for each $m$, is as in the previous section. Furthermore, we continue to model reshuffling as a reduction in the discount factor, as, given the continuum of agents and a finite set of assignments and states, players can be matched to a new team in the same reactive assignment and state.
@@
Lastly, the dynamic team structure must respect a task-assignment constraint—the equivalent of (4): for each task, the steady-state mass of teams assigned to states corresponding to that task, summed across reactive assignments, must satisfy an exogenous constraint. Formally, let $\mathcal{S}^{t}(m,\overline{m})\subseteq\mathcal{S}$ denote the set of states $s=[(t_{\tau},g_{\tau})]_{\tau=0}^{K}\in\mathcal{S}$ such that $t_{K}=t$ for $t\in\mathcal{T}$, i.e., those where the current task is task $t$. The overall constraint on a team structure is:
@@
In the assignment from Example 1, using the steady-state distribution over states given in (7), the share on task $t_{G}$ is $\frac{a_{G}}{2+x}+\frac{1-a_{G}}{x+2}=\frac{1}{x+2}$ (and consequently the share on task $t_{B}$ is $\frac{1+x}{2+x}$). Depending on the overall distribution of teams across tasks that must be satisfied and the exact value of $x$, a team structure $\lambda$ may have to put weight on other reactive assignments in order to satisfy (8).
This extension of the model, which allows the designer a richer set of assignments, again nests all previous formulations. For instance, a random, but static, task assignment as in Section 4 is captured by reactive assignment $m$ being a constant function.
The next “technical” lemma establishes that Properties 16 continue to imply a simple on-path equilibrium structure in which players only fully cooperate or fully not cooperate. This simplifies the analysis, as we do not need to consider more intricate potential structures in which some players cooperate, and others do not. Importantly, this result implies that there is a well-defined social value given any reactive team structure.
@@
$\sum_{m\in{\mathrm{supp}}(\lambda)}\lambda(m)\left(\sum_{s\in\mathcal{G}(m,r_{m})}\overline{m}(s)V_{G}+\sum_{s\in\mathcal{B}(m,r_{m})}\overline{m}(s)V_{B}\right).$ (9)
To illustrate, the social value coming from our running example of a reactive assignment with steady state given in (7) is proportional to $\frac{a_{G}}{2+x}V_{G}+\left(\frac{a_{B}}{2+x}+\frac{xa_{B}}{2+x}\right)V_{B}$, assuming there is cooperation throughout.